OpenStudy (anonymous):

How do I find the square root of a number? :) For example I was thinking 36 would be 6. Don't give me an answer to the question though, just tell me if I'm right. If I'm not then just explain and see if I can figure it out. :)

OpenStudy (anonymous):

you are right about 36

OpenStudy (anonymous):

sqrt of a number is 2 exact numbers that when multiplied equal whatever is inside the sqrt sign. sqrt(25)= 5 because 5*5 = 25

OpenStudy (anonymous):

symbolically it would be sqrt(x)=a and therefore a*a=x

OpenStudy (anonymous):

Can it be -6 also?

OpenStudy (anonymous):

yes mam

jimthompson5910 (jim_thompson5910):

timo86m you're mixing up two concepts

OpenStudy (anonymous):

-6*-6 = 36 indeed

jimthompson5910 (jim_thompson5910):

the equation x^2 = 36 has two solutions, which are x = 6 or x = -6 that is true

jimthompson5910 (jim_thompson5910):

but the square root function is a function keyword: function each input generates exactly one output

jimthompson5910 (jim_thompson5910):

the square root of 36 is just 6 (not -6 at the same time) the convention is to make the output positive or 0 if you want the negative version of it, you write \[\Large -\sqrt{36}\]

OpenStudy (anonymous):

OK, that's what I thought. I just wanted to make sure. Thanks so much! :)